4. Let X have p.d.f. fx(1),-1 < 2. Find the p.d.f. of Y-X2
Answer #1

here as we know from transformation method:

pdf of Y: f(y)=|dx/dy| *fx(g-1(y))

here X =-/+√y

|dx/dy| =1/(2√y)

hence f(y)=(1/(2√y))*(fx(√y))+fx(-√y)))

therefore for 0 <y<1

f(y)=(1/(2√y))*((2/9)*(√y+1)+(2/9)*(√y+1))

f(y)=(2/(9√y))*(√y+1)

for 1<y<4

f(y)=(1/(2√y))*((2/9)*(√y+1)+0)

f(y)=(1/(9√y))*(√y+1)

hence pdf f(y) =(2/(9√y))*(√y+1)   for 0 <y<1

       and f(y)=(1/(9√y))*(√y+1)   for 1<y<4

otherwise f(y) =0